OpenStudy (anonymous):

Evaluate the integral (8/x-1/x^5+1/3x^7) dx

OpenStudy (anonymous):

Is that \[\int\limits (\frac{ 8 }{ x } - \frac{ 1 }{ x } + \frac{ 1 }{ 3 }x^7) dx\]?

OpenStudy (anonymous):

^meant to write \[\int\limits\limits (\frac{ 8 }{ x } - \frac{ 1 }{ x ^5} + \frac{ 1 }{ 3 }x^7) dx\]

OpenStudy (anonymous):

\[1/3x ^{7}\]

OpenStudy (anonymous):

everything else is right

OpenStudy (anonymous):

Ok, so you can split this off into multiple integrals like:\[8 \int\limits (\frac{ 1 }{ x }) dx - \int\limits (\frac{ 1 }{ x^5 }) dx + \frac{ 1 }{ 3 } \int\limits (\frac{ 1 }{ x^7 }) dx\] Do you understand what I did there and how to go from there?

OpenStudy (anonymous):

how do I handle the 1/x^5

OpenStudy (anonymous):

You can treat that as \[x ^{-5}\] and just apply the power rule